OpenStudy (faman39):

Karim borrowed $120 from his older brother for the purchase of a skateboard. He promised to pay back the loan in 6 months at a simple annual interest rate of 15%. How much will Karim owe his brother at that time?

OpenStudy (anonymous):

E

OpenStudy (anonymous):

I = prt = 120 * 15% * 1/2 = $9 -> Total owned: 120 + 9 = $129